An aeroplane heads due north at 500 km/h. It experiences a 80 km/h crosswind flowing in


the direction N60^°E.



(b) Determine the speed of the aeroplane. (Leave your answer in terms of square root) (3)

1
Expert's answer
2022-04-29T13:59:10-0400
"\vec{v}=500\vec{j}"

"\vec{u}=80\sin60\degree \vec{i}+80\cos60\degree\vec{j}"

"\vec{v}_{res}=\vec{v}+\vec{u}"

"=40\sqrt{3}\vec{i}+540\vec{j}"

"|\vec{v}_{res}|=\sqrt{(40\sqrt{3})^2+(540)^2}=20\sqrt{741}(km/h)"